Evaluate: 𝑛^5 − 4𝑝^0 + (𝑝 − 𝑛)^2 − 𝑝𝑛 if 𝑛 = −2 and 𝑝 = 3
Answer by Boreal(15235)

𝑛^5 − 4𝑝^0 + (𝑝 − 𝑛)^2 − 𝑝𝑛 if 𝑛 = −2 and 𝑝 = 3
-2^5-4+(3-(-2))^2-3(-2)=
-32-4+25+6
=-5